• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Bailey

    Log entry time 07:57:14 on August 4,2005

    Entry number 150639

    keyword=Owl Shift Summary

    Owl shift summary
    Shift workers:
    S. Bailey, A. Kolarkar, Y. Prok

    ------------------------------
    23:21 Start Production Run 4096, IHWP In, 35 uA
    00:17 End Production Run 4096 - 44,080 good parity events
    00:18 Start Production Run 4097, IHWP In, 35 uA
    01:14 End Production Run 4097 - 43,734 good parity events
    01:15 Start Production Run 4098, IHWP In, 35 uA
    01:28 Cut beam because of fluctuations in CHL flow.
    01:38 End Production Run 4098 because target fluctuations
    - 4,689 good parity events
    -------------------------------
    01:48 Target to manual mode to stabilize temperature, pressure, heat load.
    01:48 Paged target expert, no response
    02:00 Paged RC
    02:27 Asked MCC to call Cryogen expert on-call
    02:33 Cryogen expert on-call Jonathan Creel said he will look at fluctuations and call us back.
    03:00 Target to PID mode to check behavior. Same irregularities. Back to manual mode.
    04:15 Jonathan reported that cryogenics looks ok. The supply loads, pressures, and recovery are all flat. He cannot find anything that matches the waveform we see in the CHL.
    04:45 Changed the fan speed from 70 Hz to 48 Hz. Returned target to PID mode to observe target behavior. Temperature, pressure, heat load still oscillates irreguarly.
    04:55 Reboot target IOC. Everything ok again. Changed fan speed to 60 Hz.
    -------------------------------
    05:04 Asked for 5 uA, checked raster size
    MCC said that I should have turned off beam modulation while we diagnosed the target problem. Other halls suffered and it wasn't necessary. I didn't know this.
    05:05 Asked for 35 uA, increased in increments of 10 uA
    05:08 Start Production Run 4099, IHWP In, 35 uA
    05:30 Start Compton Run 11705
    06:04 End Production Run 4099 - 39,960 good parity events
    06:04 Start Production Run 4100, IHWP In, 35 uA
    06:04 Increased current from 35 uA to 36 uA
    07:00 End Production Run 4100 - 46,444 good parity events
    07:00 Start Production Run 4101, IHWP In, 35 uA
    07:56 End Production Run 4101
    07:56 Start Production Run 4102, IHWP In, 35 uA